OpenStudy (anonymous):

D/-3+ 7 = 10 what does d =?

OpenStudy (anonymous):

d=40

OpenStudy (anonymous):

If it is (d/-3) + 7 = 10 then, d/-3 = 10 - 7 d = 3 * -3 = -9 d =-9

OpenStudy (anonymous):

-3+7=4 40/4=10

OpenStudy (anonymous):

is d just a variable or does it mean diameter in this equation

OpenStudy (anonymous):

whose answer was right?

OpenStudy (anonymous):

helping tutors